Percival Company wishes to sell wooden beams to home builders. The current market price of the beams is​ $950, and Percival knows it must accept the market price. ​ Currently, the beams cost Percival​ $809 to produce. The company wishes to make a profit equal to​ 16% of the price. Which of the following strategies would help Percival accomplish their​ objective?

A. Increase the volume of sales.

B. Reduce production costs by​ $11 per unit.

C. Increase the sales price to​ $961.

D. Increase the production cost by​ $22 per unit.
Answer #1

B.reduce production costs by 11dollars per unit.

reduce the production cost to 11 dollars per unit so that it is possible to earn 16%profit on sale price.

950 * 16% = 152.

production cost = 950 - 12 = 798.